a) If X~Bin(45,0.6), what is E[X]?

b) If X~Bin(45,0.4), what is Var[X]?
Answer #1

a)

E(X) = np = 45 * 0.6 = 27

b)

Var(X) = npq

= 45*0.4*0.6

= 10.8.

Please rate